A manufacturer inspects 800 batteries and finds that 796 have no defects... there are 25,000 batteries in storage. How many batteries are likely to have no defects?
Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(3719)

796/800x25000=24875 batteries are likely to have no defects.

Answer by stanbon(75887)

A manufacturer inspects 800 batteries and finds that 796 have no defects... there are 25,000 batteries in storage. How many batteries are likely to have no defects?
-----------------------
Emperical probabilty of no defect = 796/800
--------------
Expected number of no defect in 25000 = (796/800)*25000
=======================
cheers,
stan .